OpenStudy (anonymous):

Given f(x) = –5x + 1, what is f(–2)? A.3/5 B.11 C.-9 D.1

OpenStudy (anonymous):

Just plug in -2 for every x you see in the function. f(-2) = -5(-2) + 1 f(-2) = 10 + 1 f(-2) = 11

OpenStudy (anonymous):

Do you understand what I did?

OpenStudy (anonymous):

Feel free to message me if you don't and I'll clarify.